In the past week and this week the contractors on site continue to work on their projects. Veliz Construction is half way done with the application of the silver tape to lower the emissivity of the control and services building next to the dome. We expect them to take about 2 weeks to finish the rest of the tape work if the afternoon rains keep their current pattern and stop their work. ARC Mechanical continue to populate the vault with freon, glycol and other lines as well as connecting up to the inside of the building. This work progresses slowly.
We have gotten the approval for the opto-mechanical technician position. If you have interest or know someone interested in working in the beautiful Davis Mountains and working with the optics of one of the largest telescopes in the world please have them apply to: job number 140715014211
Last week we had the metrology team out from Austin and reinstalled the VAT, tip-tilt camera and VAT on the telescope. This trip we had much greater success. We were able to get the system aligned and got a return from the DMI. We also were able to turn the VAT around and point at the sky. We got our first trajectory and followed the star we actually desired for 20 minutes in which time the star did not drift appreciably. We did have some problems getting trajectories to start so that was the focus of the work this week. After a few days of digging we found a variable that was defined in the wrong part of the code and with the variable moved to the right location we were able to run 8 trajectories one after another and also ran trajectories that simulated stars moving in the East and North (we had been working in the South were there is little rotation and the trajectories move more quickly). This is a big break through!